OpenStudy (anonymous):

5x-1 over 4 > -2 (x+5) ---- I need help understanding this problem.

OpenStudy (the_fizicx99):

\(\ \dfrac{5x - 1}{4} > -2(x + 5) \) ?

OpenStudy (the_fizicx99):

I see your screenshot now, lol Start by multiplying both sides by 4. By doing this you're eliminating the 4 dividing 5x - 1. Damn... I can't use the Draw option :/ \(\ \sf 4\dfrac{5x - 1}{4} \longrightarrow 5x - 1 \) What you do to one side, you \(\ must\) do it to the other to keep the equation balanced. Therefore, 5x - 1 = 4(-2x - 10); I already distributed the -2 in -2(x + 5)

OpenStudy (the_fizicx99):

Quick question, do you speak anything other than English?

OpenStudy (anonymous):

No. Why ?

OpenStudy (the_fizicx99):

Your grammar seems different from what I'm used to speaking to natives. Lol

OpenStudy (snowsurf):

Did this math problem ask for interval solution that would make this inequality true?

OpenStudy (the_fizicx99):

I just noticed a typo, I'm used to using "=," sorry. 5x - 1 > 4(-2x - 10) After solving you'd get x >-3

OpenStudy (anonymous):

Thank you! The work you provided me was correct. You helped me a lot with this problem and many others.